OpenStudy (anonymous):

\[\sqrt[3]{5/2}\]. rationalize the denominator

OpenStudy (anonymous):

i git an index of 3 with radical 20) over 2

OpenStudy (anonymous):

\[\sqrt[3]{20}/2\]

OpenStudy (anonymous):

can anybody confirm?

OpenStudy (anonymous):

that is correct i believe

OpenStudy (anonymous):

you can simplify more if you want

OpenStudy (anonymous):

woot, thanks!

OpenStudy (anonymous):

actually no you can't simplify more lol that is correct

OpenStudy (anonymous):

cube root of (5/2) = cube root of (20/8) = cube root of 20/ cube root of 8 = \[\sqrt[3]{20} \div 2\]

OpenStudy (anonymous):

that's proof